OpenStudy (ashleyisakitty):

Given the proportion a/b=14/60, which ratio completes the equivalent proportion a/14?

OpenStudy (anonymous):

this one is easy

OpenStudy (ashleyisakitty):

will you explain?

OpenStudy (anonymous):

do u want the answer in the explanation

OpenStudy (anonymous):

or how to come up with it

OpenStudy (ashleyisakitty):

the explanation to the answer would be best

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what type of math is this

OpenStudy (ashleyisakitty):

geometry

OpenStudy (anonymous):

okay

OpenStudy (campbell_st):

well if you have \[\frac{a}{b} = \frac{14}{60}\] divide both sides of the equation by 14 which gives \[\frac{a}{14b} = \frac{1}{60}\] next multiply both sides of the equation by b to get your answer

OpenStudy (anonymous):

do u see the Absolute Value a/b=14/60

OpenStudy (anonymous):

im a little of on my geometry

OpenStudy (ashleyisakitty):

@campbell_st, latex is broken

OpenStudy (ashleyisakitty):

the absolute value of what exactly? x?

OpenStudy (campbell_st):

ok... so you have a/b = 14/60 divide both sides of the equation by 14 gives a/(14b) = 1/60 all you need to do is multiply boths sides of the equation by b to get the answer

OpenStudy (campbell_st):

@Ashleyisakitty thanks for the info on Latex... explains why people aren't replying to my help... lol

OpenStudy (anonymous):

by b??

OpenStudy (ashleyisakitty):

multiply both sides by 60?

OpenStudy (anonymous):

I was trying to find it out but Im weak on my geometry

OpenStudy (campbell_st):

ok... so you have a/(14b) = 1/60 multiply both sides of the equation by b... that removes it from the denominator of the 1st ratio.. does that make sense..?

OpenStudy (ashleyisakitty):

Is the answer a/60?

OpenStudy (anonymous):

YESSSS

OpenStudy (campbell_st):

nope...

OpenStudy (anonymous):

the way I enterd it was divid the most high number

OpenStudy (anonymous):

but looks like i did it wrong

OpenStudy (campbell_st):

a/(14b) * b = 1/60 * b which gives a/14 = b/60

OpenStudy (anonymous):

comes out to be 14

OpenStudy (campbell_st):

they question is asking what's equivalent to a/14

OpenStudy (anonymous):

ohhh equivalent to a /14

OpenStudy (ashleyisakitty):

Okay now I understand what you were saying @campbell_st. Thank you.